CONFORMATIONS OF p-METHOXYPHENETHYLAMINE, A NEUROTRANSMITTER

KARISSA G. ATTICKS, \underline{ROBERT K. BOHN}, Department of Chemistry, University of Connecticut, Storrs, CT 06269-3060; YOSHI KAWASHIMA, and RICHARD D. SUENRAM, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899.

p-Methoxyphenethylamine (CH_{3}OC_{6}H_{4}C_{a}H_{2}C_{b}H_{2}NH_{2}) is a neurotransmitter. Seven conformers have been identified by electronic spectroscopy^{a} and their conformational identities characterized by analogy to simpler related compounds, empirical force field and quantum chemical calculations, and vibrational spectra. With the C_{a}-C_{b} bond lying in a plane nominally orthogonal to the benzene ring, there are nine possible conformers. With C(phenyl)-C_{a}-C_{b}-N gauche, the NH_{2} group has 3 possible orientations. These are doubled by the two possible planar orientations of the methoxy group. With C(phenyl)-C_{a}-C_{b}-N anti, the NH_{2} group again has 3 distinguishable orientations. A pair of the C(phenyl)-C_{a}-C_{b}-N gauche conformers with the N lone pair directed toward the benzene ring were deemed to be missing in the electronic spectroscopic studies. The rotational spectrum was investigated at a rotational temperature near 1 K in a pulsed-jet Fourier transform microwave spectrometer. Rotational spectra of four conformers have been assigned with preliminary rotational constant values of 2832.545, 538.362, and 509.172 MHz; 2740.111, 542.176, and 507.522 MHz; 2789.609, 535.220, and 509.950 MHz; and 2750.592, 548.056, and 507.537 MHz. They all correspond to conformers with a C(phenyl)-C_{a}-C_{b}-N gauche conformation coupled with the methoxy group ”anti” or ”syn” to the gauche NH_{2} group. Structural characterization is continuing. More lines remain to be assigned.

aS. J. Martinez III, et al., J. Molec. Spectrosc., 158, 82(1993) and I. Unamuno, et al., J. Phys. Chem. A, 104, 4364(2000).